# Beyond Sequence: Impact of Geometric Context for RNA Property Prediction

### v1

@@ 1 Introduction
Figure 1: Overview of the study. (a) Left panel: RNA sequences represented in 1D, 2D, and 3D structures, processed by 1D sequence, 2D GNN, and 3D GNN models. Our analysis includes prediction error, robustness and generalization to sequencing noise, and performance under limited training data and partial labels. (b) Right panel: Comparative performance of 1D, 2D, and 3D methods across experimental conditions. Histograms show RMSE performance, relative RMSE changes with increasing noise, and data requirements for optimal performance. Lower values indicate better performance in all metrics.
@@
RNA plays a central role in the machinery of life, serving as a crucial intermediary between nucleotide and amino acid worlds [39]. Beyond its messenger role, RNA is involved in diverse biological processes, including gene regulation, catalytic activity, and structural support within ribosomes [44, 50]. This versatility makes RNA a key target for fundamental biological research and therapeutic interventions. As our understanding of RNA complexity grows, so does the need for advanced computational tools for its analysis.
@@
Modeling RNA is challenging due to its intricate secondary and tertiary structures, dynamic conformational changes, and interactions with cellular components [19]. Furthermore, RNA analysis is hindered by the practical challenges of RNA data acquisition which include sequencing errors [33], batch effects [53], incomplete sequencing [2], partial labeling [60], and high costs of obtaining large labeled datasets [7]. Moreover, RNA molecules can be represented in different ways: as a 1D nucleotide sequence, a 2D graph of base pairings, or a 3D atomic structure. Each representation highlights different aspects of RNA, presenting both opportunities and challenges for model design and selection (Fig. 1).
@@
In this work, we systematically study the performance of various machine learning models for RNA property prediction, extending beyond traditional sequence-based approaches [20, 49, 15] to include methods that process RNA with its 2D or 3D geometry. While 2D and 3D RNA representations offer potentially richer information, they also present unique challenges. In the absence of high-quality experimental data, accessing 2D or 3D RNA structures requires running structure prediction algorithms prone to noise and mistakes, especially in the presence of sequencing errors [41, 57]. A few mutations in the nucleotide sequence owing to sequencing mistakes can hugely alter the 2D and 3D structure (Fig. 4), potentially undermining the benefit of using additional geometric context. Furthermore, real-world RNA datasets often suffer from partial labeling [60] and scarcity of training samples [62], which may affect geometry-aware methods differently than sequence-based approaches. These challenges raise a fundamental question: to what extent does explicit 2D and 3D geometry contribute to RNA property prediction, and under what circumstances might it offer advantages over geometry-free sequence models?
@@
Our study reveals several key insights: (i) 2D models generally outperform 1D models, with spectral GNNs reducing prediction error by about 12% on average across all datasets, highlighting the importance of explicitly considering RNA structural information; (ii) 3D equivariant GNNs outperform 1D and some 2D methods in the noise-free scenario but are sensitive to noise, exhibiting up to a 56% decrease in prediction quality under high sequencing error rates; (iii) geometry-free sequence models remain the most robust to sequencing noise, showing only a 14-27% increase in prediction error compared to noise-free conditions, however, they require around 25×2-5$\times$2 - 5 × more training data to match the performance of geometry-aware models.
@@ 2.1 Datasets
The datasets vary in size based on the number of sequences and sequence lengths: the small dataset Tc-Riboswitches [17], the medium datasets Open Vaccine COVID-19 [60] and Ribonanza-2k [22], and the large dataset Fungal [62]. All datasets provide regression labels. Detailed statistics for these datasets are provided in Appendix C.1.
@@ 2.2 Data Preprocessing and Curation
For the OpenVaccine COVID-19 dataset, we filter out sequences with a signal-to-noise ratio (SNR) below 1, as recommended by the dataset authors [60], to ensure that only sequences with a significant signal relative to background noise are included, thereby enhancing the reliability of modeling. For the other datasets, we use the original sequences since no SNR annotations are available.
Since all the RNA datasets come with sequences only, we employ EternaFold [59] and RhoFold [45] to infer 2D and 3D molecular structures respectively. We selected EternaFold and RhoFold due to their state-of-the-art performances acknowledged in recent works [58, 60, 22] Additionally, RhoFold typically runs in seconds to a minute per sequence, unlike other 3D structure prediction tools which usually take hours, and hence not suitable for large datasets.
@@ 2.3 Models
We select well-established model architectures recognized for their state-of-the-art performance for molecular property prediction tasks in various domains. 1D Model: Transformer1D [24, 20]; 2D Models: GCN [27, 61], GAT [55, 67], ChebNet [11, 28], Transformer1D2D [21], Graph Transformer [47, 31], and GraphGPS [38, 70]; 3D Models: SchNet [42, 18] and EGNN [40].

### v2

@@ 1 Introduction
Figure 1: Overview of the study. (a) Left panel: RNA sequences represented in 1D, 2D, and 3D structures, processed by 1D sequence, 2D GNN, and 3D GNN models. Our analysis includes prediction error, robustness and generalization to sequencing noise, and performance under limited training data and partial labelings. (b) Right panel: Comparative performance of 1D, 2D, and 3D methods across experimental conditions. Histograms show RMSE performance, relative RMSE changes with increasing noise, and data requirements for optimal performance. Lower values indicate better performance in all metrics.
@@
RNA plays a central role in the machinery of life, serving as a crucial intermediary between nucleotide and amino acid worlds (Sahin et al., 2014). Beyond its messenger role, RNA is involved in diverse biological processes, including gene regulation, catalytic activity, and structural support within ribosomes (Sharp, 2009; Strobel et al., 2016). This versatility makes RNA a key target for fundamental biological research and therapeutic interventions. As our understanding of RNA complexity grows, so does the need for advanced computational tools for its analysis.
@@
Modeling RNA is challenging due to its intricate secondary and tertiary structures, dynamic conformational changes, and interactions with cellular components (Han et al., 2015). Furthermore, RNA analysis is hindered by the practical challenges of RNA data acquisition which include sequencing errors (Ozsolak & Milos, 2011), batch effects (Tran et al., 2020), incomplete sequencing (Alfonzo et al., 2021), partial labeling (Wayment-Steele et al., 2022b), and high costs of obtaining large labeled datasets (Byron et al., 2016). Moreover, RNA molecule can be represented in different ways: as a 1D nucleotide sequence, a 2D graph of base pairings, or a 3D atomic structure. Each representation highlights different aspects of RNA, presenting both opportunities and challenges for model design and selection (Fig. 1).
@@
In this work, we systematically study the performance of various machine learning models for RNA property prediction, extending beyond traditional sequence-based approaches (He et al., 2021; Soylu & Sefer, 2023; Prakash et al., 2024; Yazdani-Jahromi et al., 2025) to include methods that process RNA with its 2D or 3D geometry. While 2D and 3D RNA representations offer potentially richer information, they also present unique challenges. In the absence of high-quality experimental data, accessing 2D or 3D RNA structures requires running structure prediction algorithms prone to noise and mistakes, especially in the presence of sequencing errors (Schneider et al., 2023; Wang et al., 2023). A few mutations in the nucleotide sequence owing to sequencing mistakes can hugely alter the 2D and 3D structure (Fig. 4), potentially undermining the benefit of using additional geometric context. Furthermore, real-world RNA datasets often suffer from partial labeling (Wayment-Steele et al., 2022b) and scarcity of training samples (Wint et al., 2022), which may affect geometry-aware methods differently than sequence-based approaches. These challenges raise a fundamental question: to what extent does explicit 2D and 3D geometry contribute to RNA property prediction, and under what circumstances might it offer advantages over geometry-free sequence models?
@@
Our study reveals several key insights: (i) 2D models generally outperform 1D models, with spectral GNNs reducing prediction error by about 12% on average across all datasets, highlighting the importance of explicitly considering RNA structural information; (ii) 3D equivariant GNNs outperform 1D and some 2D methods in noise-free scenario but are sensitive to noise, exhibiting up to a 56% decrease in prediction quality under high sequencing error rates; (iii) geometry-free sequence models remain the most robust to sequencing noise, showing only a 14-27% increase in prediction error compared to noise-free conditions, however they require around 25×2-5$\times$2 - 5 × more training data to match the performance of geometry-aware models.
@@ 2.1 Datasets
The datasets vary in size based on the number of sequences and sequence lengths: the small dataset Tc-Riboswitches (Groher et al., 2018), the medium datasets Open Vaccine COVID-19 (Wayment-Steele et al., 2022b) and Ribonanza-2k (He et al., 2024), and the large dataset Fungal (Wint et al., 2022). All datasets provide regression labels. Detailed statistics for these datasets are provided in Appendix F.1.
@@ 2.2 Data Preprocessing and Curation
For the OpenVaccine COVID-19 dataset, we filter out sequences with a signal-to-noise ratio (SNR) below 1, as recommended by the dataset authors (Wayment-Steele et al., 2022b), to ensure that only sequences with a significant signal relative to background noise are included, thereby enhancing the reliability of modeling. For the other datasets, we use the original sequences since no SNR annotations are available.
Since all the RNA datasets come with sequences only, we employ EternaFold (Wayment-Steele et al., 2022a) and RhoFold (Shen et al., 2022) to infer 2D and 3D molecular structures respectively. We selected EternaFold and RhoFold due to their state-of-the-art performances acknowledged in recent works (Wayment-Steele et al., 2020; 2022b; He et al., 2024) Additionally, RhoFold typically runs in seconds to a minute per sequence, unlike other 3D structure prediction tools which usually take hours, and hence not suitable for large datasets.
@@ 2.3 Models
We select well-established model architectures recognized for their state-of-the-art performance for molecular property prediction tasks in various domains. 1D Model: Transformer1D (Honda et al., 2019; He et al., 2021); RNA-FM (Chen et al., 2022), SpliceBERT (Chen et al., 2023) 2D Models: GCN (Kipf & Welling, 2017; Wieder et al., 2020), GAT (Veličković et al., 2018; Ye et al., 2022), ChebNet (Defferrard et al., 2016; Knyazev et al., 2018), Transformer1D2D (He et al., 2023), Graph Transformer (Shi et al., 2020; Li et al., 2022), and GraphGPS (Rampášek et al., 2022; Zhu et al., 2023); 3D Models: SchNet (Schütt et al., 2017; Han et al., 2022), EGNN (Satorras et al., 2021), FAENet (Duval et al., 2023), DimeNet (Gasteiger et al., 2020), GVP (Jing et al., 2020), and FastEGNN (Zhang et al., 2024). Detailed descriptions of all the models can be found in Appendix Sec. B.
Training and evaluation
All models were trained on a NVIDIA A100 GPU. To ensure hyperparameter parity for each baseline, hyperparameters were optimized using Optuna (Akiba et al., 2019), restricting the search to models with fewer than 10 million parameters that fit within the GPU memory constraint of 80GB. All model hyperparameters, training, and evaluation details are reported in Appendix H. We ran all models for $5$ random data splits (train:val:test split of 70:15:15) and we report average performance with a standard deviation across splits. The mean column-wise root mean squared error (MCRMSE), introduced in Wayment-Steele et al. (2020), is used as the evaluation metric. It is defined as $\text{MCRMSE}(f,D)=\sqrt{\frac{1}{n}\sum_{i=1}^{n}\left(\hat{y}_{i}-y_{i}\right)^{2}}$, where $f$ represents a model, $D$ is the dataset, and $\hat{y}_{i}$ and $y_{i}$ are the predicted and true values for data point $i$.
3 Task Definitions
Task 1: Impact of structural information on prediction performance
Task 2: Model efficiency in limited training data settings
Acquiring high-quality comprehensive RNA datasets is often challenging and resource-intensive thus limiting the amount of labeled data for training (Teufel & Sobetzko, 2022; Byron et al., 2016). This task aims to investigate how model performance depends on the amount of training data used, evaluating the sample efficiency of each family of models. In other words, given a dataset $D=\{X,Y\}$, let $D_{\alpha}=\{X_{\alpha},Y_{\alpha}\}$ be a subset where the training set is reduced to a fraction $\alpha$. We train models on different sets of $D_{\alpha}$ datasets with decreasing $\alpha$.
Task 3: Performance with partial sequence labeling
Due to the high cost of measuring properties for every nucleotide in RNA sequence, real-world datasets often contain partial annotations (Wayment-Steele et al., 2022b) where labels are only available for the first small part of the sequence. This task is relevant for nucleotide-level datasets and it aims to investigate how well a model can generalize to a whole RNA sequence when labels are only available for a portion of it.
Task 4: Robustness to sequencing noise
Acquiring RNA data requires sequencing. In practice sequencing procedure may introduce sequencing errors (random mutations of nucleotides) that vary depending on the sequencing technology and platform (Ozsolak & Milos, 2011; Fox et al., 2014). These errors affect the raw sequence data, and propagate to structural noise in 2D and 3D. The goal of this task is to assess how well models can maintain reliable performance when trained and tested under the same distribution of realistic levels of sequencing noise observed in practice, ensuring robustness across a consistent noise environment. This reflects real-world cases where a specific sequencing method produces noisy data, but the noise characteristics are stable across training and deployment.
Task 5: Generalization to Out-of-Distribution (OOD) data
This task focuses on a different practical challenge: models trained on high-quality RNA sequences are often deployed in conditions where the data exhibits different noise characteristics due to batch effect (Tran et al., 2020) or the use of different sequencing platforms (Tom et al., 2017). Here, the objective is to evaluate how well models generalize to OOD datasets with different levels of sequencing noise, assessing the extent of performance degradation as noise levels increase. This task simulates the scenario where a model encounters noisier data than it was trained on, highlighting its ability to adapt to unexpected experimental conditions.
For detailed descriptions and motivations of the five task settings, please refer to Appendix D.
4 Experiments and Results
4.1 Impact of explicit geometry learning on model performance
We begin by addressing Task 1, where we compare the performance of model families when trained and evaluated on the downstream RNA datasets. Additionally, we provide runtime and memory comparison in Appendix C.
Model COVID Ribonanza Tc-Ribo Fungal
1D model
RNA-FM 0.591$\pm$0.081 0.990$\pm$0.144 0.693$\pm$0.001 1.420$\pm$0.028
SpliceBERT 0.588$\pm$0.077 1.022$\pm$0.144 0.708$\pm$0.003 $1.435\pm$0.059
2D model
Transformer1D2D 0.305$\pm$0.012 0.514$\pm$0.004 0.633$\pm$0.001 OOM
GCN 0.359$\pm$0.009 0.595$\pm$0.006 0.701$\pm$0.004 1.192$\pm$0.077
GAT 0.315$\pm$0.006 0.534$\pm$0.006 0.685$\pm$0.024 1.112$\pm$0.035
ChebNet 0.279$\pm$0.007 0.468$\pm$0.002 0.621$\pm$0.022 0.973$\pm$0.003
3D model (without pooling)
EGNN 0.480$\pm$0.025 0.808$\pm$0.023 0.725$\pm$0.002 OOM
SchNet 0.499$\pm$0.003 0.843$\pm$0.004 0.696$\pm$0.008 OOM
FAENet 0.486$\pm$0.010 0.834$\pm$0.003 0.703$\pm$0.011 OOM
DimeNet 0.497$\pm$0.012 0.855$\pm$0.006 0.712$\pm$0.004 OOM
GVP 0.467$\pm$0.010 0.797$\pm$0.012 0.744$\pm$0.004 OOM
FastEGNN 0.477$\pm$0.005 0.816$\pm$0.014 0.753$\pm$0.001 OOM
3D model (with nucleotide pooling)
EGNN (pooling) 0.364$\pm$0.003 0.619$\pm$0.007 0.663$\pm$0.010 OOM
SchNet (pooling) 0.390$\pm$0.006 0.685$\pm$0.006 0.655$\pm$0.038 OOM
FastEGNN (pooling) 0.444$\pm$0.003 0.753$\pm$0.015 0.710$\pm$0.011 OOM
2D models consistently outperform 1D model
Results in Table 1 reveal that 2D methods consistently outperform the 1D sequence model across all datasets. Notably, the Transformer1D2D model, which simply augments the attention matrix with adjacency features alongside, achieves around 10% lower prediction MCRMSE on average across datasets than its geometry-free counterpart. This suggests that explicitly incorporating structural information is crucial, as learning from sequence data alone proves to be insufficient. Further experiments, detailed in the Appendix F.4, investigate the learned attention maps of both the Transformer1D2D and the Transformer1D model and their correlation with structural information and reveal that Transformer1D2D attention maps are much more closely aligned with the topological structure of nucleotide graph, reinforcing the conclusion that explicit encoding of structural information is essential for improved performance.
For the foundation models, RNA-FM and SpliceBERT, we observe that for 2 out of 4 datasets (Covid and Ribonanza-2k), RNA foundation models perform worse than the simple supervised transformer baseline whereas for the other two datasets (Tc-Riboswitches and Fungal), transformer and RNA foundation models achieve similar performance. This is consistent with recent works in multiple biology-related domains demonstrating specialized foundation models are yet to surpass simple supervised learning baselines (Xu et al., 2024b; Kedzierska et al., 2023; Yang et al., 2024). Hence, we choose Transformer1D as the model of choice for subsequent experiments.
Spectral GNN outperforms spatial GNNs in 2D
ChebNet, a spectral method, outperforms spatial methods such as GCN, GAT, Graph Transformer, and GraphGPS, achieving a prediction MCRMSE 2.5% lower than the next best 2D model across datasets. Spatial GNNs aggregate node features layer by layer, emphasizing local information within a fixed distance. While computationally efficient, these methods are limited by the 1-Weisfeiler-Lehman (WL) test, which constrains the expressive power of node-based updates (Xu et al., 2018). In addition, spatial GNNs may suffer from a limited receptive field while spectral methods approximate global graph features, enabling a global receptive field since the first layer Wang & Zhang (2022); Bo et al. (2023); Xu et al. (2024a). This allows ChebNet to effectively process global information which is important for RNA data due to potential long-range interaction between nucleotides.
Challenges of modeling geometric context in all-atom resolution
Contrary to our expectations, 3D models at all-atom resolution (EGNN w/o pooling, SchNet w/o pooling DimeNet (w/o pooling), FAENet (w/o pooling), GVP (w/o pooling) and FastEGNN (w/o pooling) in Table 1) show relatively high prediction MCRMSE across datasets, underperforming compared to 1D and 2D methods. We hypothesize this is due to two factors.
First, all-atom 3D models rely on a limited local neighborhood of adjacent atoms, limiting their receptive fields and preventing them from capturing long-range dependencies (see Appendix Table F.1), which can be crucial for determining RNA properties (Shetty et al., 2013; Alshareedah et al., 2019). Expanding the local neighborhood for these methods becomes challenging due to the overwhelming scale of large molecular systems in all-atom resolution. Second, the performance of 3D models is often limited by the inherent inaccuracies in 3D structure prediction tools, which are generally less reliable compared to 2D structure prediction methods (Ponce-Salvatierra et al., 2019). While FastEGNN is designed for larger molecules, its reliance on the center of mass virtual node initialization may not align well with tasks like RNA property prediction, which inherently have a sequence prior. This may explain the limited performance of FastEGNN in this context and RNA-specific virtual node initialization and carefully designed pooling mechanisms may be needed to further adopt this architecture for large-molecules with a sequence prior.
To address the receptive field limitation of all-atom methods, we employ a biological prior by pooling atomic features into nucleotide-level representations after a few layers of all-atom operations. This strategy is aligned with RNA’s natural secondary structure, where atoms group into nucleotides, and nucleotides form the complete RNA molecule (Deng et al., 2023). This novel strategy allows us to maintain all-atom resolution in the initial layers while increasing the receptive field via pooling. By balancing the number of all-atom and nucleotide layers as hyperparameters, we can balance fine- and coarse-grained all-atom and nucleotide resolution. Compared to other 3D models, EGNN, FastEGNN, and SchNet perform slightly better, and hence we introduce the nucleotide pooled variants for these models. We call 3D models with pooling EGNN (nuc. pooling), FastEGNN (nuc. pooling), and SchNet (nuc. pooling). This strategy significantly enhances 3D model performance, reducing prediction MCRMSE by $\sim$10compared to the original EGNN, SchNet, and FastEGNN and outperforming 1D models on the Ribonanza-2k and Tc-Riboswitches datasets by 5% on average. On the COVID dataset, EGNN (nuc. pooling) matches the Transformer1D model but still trails behind 2D models. All subsequent experiments report results using nucleotide-pooled versions of the 3D models.
Next, we investigated the second hypothesis regarding higher noise in 3D structures by quantifying variability in predicted structures across different 3D prediction tools in Appendix G. We observe substantial variability (between 11-45Å RMSD across structures given by different 3D structure prediction tools), suggesting considerable noise in 3D predictions, which likely contributes to the poorer performance of 3D models.
We also evaluate the impact of partial property labels for nucleotide-level tasks, a common occurrence owing to costly experimental measurements (Wayment-Steele et al., 2020) to identify which models are best suited to handle the challenges of incomplete labels in RNA property prediction. For this, we use the COVID and Ribonanza datasets as these datasets contain nucleotide-level labels. We train the models using all training data but with varying proportions of labeled nucleotides (20%, 40%, 60%, 80%, and 100%) per sequence, thus simulating incomplete or sparse labeling, while testing on fully labeled test sets (Appendix Fig. 8(b) for illustration).
(a) COVID
(b) Ribonanza
(c) Fungal
Figure 2: Performance vs. fraction of training data across various datasets. Model performance improves with increasing data, with lower MCRMSE across all models. 2D models consistently outperform 1D models, particularly in low-data regimes, underscoring the value of structural information for generalization. Dotted, solid, and dashed lines denote 1D, 2D, and 3D methods, respectively, which applies consistently throughout all figures in this paper.
(a) COVID
(b) Ribonanza
More training data improves performance
Unsurprisingly, across all models and datasets, a clear trend emerges: increasing the amount of available data, whether through higher number of training data points or greater proportion of available labels leads to improved performance (Fig. 2, Fig. 3, Appendix Tables in Sec. I.1I.2). However, the degree of performance improvement varies significantly between model types, as analyzed next.
2D models excel in low data and partial label regimes
Noise Ratio
0 0.1 0.2 0.3
1D Refer to caption Refer to caption Refer to caption Refer to caption
2D Refer to caption Refer to caption Refer to caption Refer to caption
3D Refer to caption Refer to caption Refer to caption Refer to caption
4.3 Model Robustness and Generalization Under Sequencing Noise
Tasks 4 and 5 both deal with model performance with noise in the data, but focus on different aspects, robustness to noise, and ability to generalize across unseen noise distributions. As explained in Sec. 3, sequencing noise is common depending on the sequencing method and platform used (Fox et al., 2014), thereby introducing errors in sequences that propagate into 2D and 3D structures. Additionally one of the deployment scenarios involves models trained on high-quality clean data applied for datasets acquired under noisy conditions owing to different sequencing platforms or experimental batch effects (Tom et al., 2017).
  • Robustness: We introduce sequencing noise into the training, validation, and test sequences to simulate realistic sequencing errors. Nucleotide mutations are applied with probabilities {0.05, 0.1, 0.15, 0.2, 0.25, 0.3}, mirroring typical sequencing error rates (Pfeiffer et al., 2018) which also propagates to the 2D and 3D structures (Fig. 4) via structure prediction tools. Importantly, these mutations are not random; the likelihood of a particular nucleotide mutating into another varies, as is well documented in sequencing studies Pfeiffer et al. (2018). Our noise model reflects these real-world mutation profiles. Crucially, while the input training, validation, and test sequences contain noise, the property labels remain clean. This again reflects practical scenarios where labels are experimentally determined independent of sequencing and thus unaffected by sequencing errors.
(a) COVID
(b) Ribonanza
(c) Fungal
(d) Tc-Riboswitches
(a) COVID
(b) Ribonanza
(c) Fungal
(d) Tc-Riboswitches
Expectedly, across both tasks, increased noise levels generally leads to worse test MCRMSE for all models, indicating a decline in prediction performance (Fig. 5, Fig. 6 and Appendix Tables in Sec. I.3I.4).
Among all models, Transformer1D demonstrates the highest robustness and generalization, exhibiting the least performance degradation as noise levels increase. Notably, in generalization experiments, Transformer1D achieves the best prediction MCRMSE on the COVID, Ribonanza, and Tc-Riboswitches datasets under higher noise levels (Fig. 6 and Appendix I.4). The reliance of Transformer1D on sequence-only information without considering geometric context, while being a weakness in other scenarios, becomes a strength in case of noisy sequences as minor sequencing errors may severely alter the predicted RNA structures (Fig. 4). While its performance on the Fungal dataset is worse overall, Transformer1D maintains remarkably consistent performance as noise increases. Transformer1D2D ranks just behind Transformer1D outperforming other 2D and 3D models. This can be attributed to Transformer1D2D’s ability to selectively focus on sequence information for noisy data rather than relying on structural data alone as the self-attention is only weakly conditioned on the graph topology.
More elaborated 2D and 3D models, which rigidly rely on structural information, are significantly more affected by noise, underperforming compared to plain sequence baseline in both robustness and generalization experiments for high noise levels. For low-to-moderate noise levels (5-10% noise), 2D methods such as ChebNet still perform the best. Interestingly, ChebNet shows the worst generalization among 2D models, although it performs on par with other methods in robustness experiments. This suggests that while ChebNet struggles with OOD noise, its performance gets better when it is also trained on the same noise level used during testing, highlighting the need for retraining for different experimental data batches/noise levels in real-world applications. Across both experiments, 3D models have poor performance for more noisy conditions, particularly with the COVID and Ribonanza datasets, due to their dependence on 3D structures which are also sensitive to the propagation of sequencing errors (Fig.4 bottom). In the smaller Tc-Riboswitches dataset, model performances vary more, likely due to limited data size, but transformer models still consistently demonstrate greater robustness to noise.
5 Conclusion
We present the first comprehensive study on the benefits and challenges of the effect of geometric context for RNA property prediction models. With providing a curated set of RNA datasets with annotated 2D and 3D structures, we systematically evaluate the performance of 1D, 2D, and 3D models under various real-world conditions, such as limited data, partial labeling, sequencing errors and out-of-distribution generalization. Our results reveal that 2D models outperform 1D and 3D models, with spectral graph neural networks excelling even in low-data and partial labeling scenarios. For 3D models, we find that their potential benefits are hindered by the limited receptive field, computational complexity, and structural noise from RNA structure prediction tools. At the same time, 1D models demonstrate better robustness compared to 2D and 3D models in noisy and OOD conditions. This study highlights the value and limitations of using geometric context for RNA modeling. Future work could focus on ensembling 1D, 2D, and 3D models for complementary strengths, and on improving 2D and 3D models to better handle noise from structure prediction tools as elaborated in Appendix E.2. Another promising direction can be to investigate advanced 3D model architectures which incorporate high-degree steerable features as discussed in Appendix E.1.
Ethics statement
This work aims to advance the computational prediction of RNA properties through the development and evaluation of machine learning models utilizing diverse representations of RNA, including 1D sequences, 2D structures, and 3D geometries. In conducting this research, we are committed to upholding high ethical standards in all aspects of our study, ensuring that our work promotes scientific integrity, transparency, and responsible use of technology.
  • Data Integrity and Fair Use: All RNA datasets used in this study were acquired from publicly available and ethically sourced repositories. Where applicable, the original data sources have been duly cited, and all efforts have been made to ensure data privacy and compliance with relevant data-sharing agreements. Our curated datasets have been handled responsibly, with proper annotations to minimize errors and misinterpretations.
  • Minimization of Bias: We recognize that RNA data, particularly when it involves limited or incomplete datasets, can introduce biases in model predictions. To mitigate this, we employ a diverse set of RNA data and systematically analyze model performance under various conditions, including label scarcity and sequencing errors. By doing so, we aim to provide balanced insights into the potential advantages and limitations of different modeling approaches, fostering responsible model selection and deployment.
  • Responsible Use of AI in Biology: The machine learning techniques used in this study are intended to assist in scientific discovery and biological understanding, with potential applications in therapeutic interventions. However, we acknowledge the importance of caution in applying computational models in sensitive domains such as healthcare. While our models are designed to improve RNA property prediction, we emphasize that these predictions should not be used in isolation for clinical or therapeutic decision-making without further validation and consideration of ethical implications.
Reproducibility Statement
We are committed to the transparency and reproducibility of our findings. All methodologies, datasets, and benchmarking environments will be made publicly available to the research community, allowing others to reproduce, verify, and extend our work. This open approach aims to advance the field of RNA research and promote collaborative progress. To ensure reproducibility, we also provide detailed descriptions of model architecture and training procedures in the main text. The appendix H contains full details on data preprocessing, hyperparameter searching and additional experimental information. All data and code will be provided to upon acceptance.
Appendix Overview
  • Appendix A: Related Work.
  • Appendix B: Models Overview.
  • Appendix C: Memory and Computational Constraints.
  • Appendix D: Detailed task descriptions.
  • Appendix E: Discussion on Advanced Model Architectures.
  • Appendix F: Additional Experimental Information.
  • Appendix G: Analysis of Noise in 3D Structures.
  • Appendix H: Reproduction.
  • Appendix I: Additional Results.
Appendix A Related Work
RNA property prediction
RNA-specific models remain scarce, likely due to limited specialized datasets. Recent advancements in sequence modeling have shown promise, particularly with foundation models like RNA-FM (Chen et al., 2022), UTRBERT (Yang et al., 2023), and RINALMO (Penić et al., 2024), which use transformer-based architectures pre-trained on large RNA sequence corpora to predict various RNA functions and structures. While RNNs and CNNs have been applied to tasks like RNA methylation and protein binding (Wang et al., 2022), they struggle with long-range dependencies (Moskalev et al., 2024). Hybrid models like RNAdegformer (He et al., 2023), combining convolutional layers with self-attention, improve predictions by capturing both local and global dependencies. Although some efforts integrate 2D structures with transformers, explicit 2D and 3D geometric modeling for RNA remains underexplored, with graph-based models mainly focusing on RNA-protein and RNA-drug interaction tasks rather than property prediction (Krishnan et al., 2024; Yan et al., 2020; Arora & Sanguinetti, 2022; Zheng et al., 2022).
RNA structure prediction
RNA 2D structure prediction has progressed from dynamic programming methods like Vienna RNAfold (Hofacker et al., 1994) to deep learning-based tools like SPOT-RNA2 (Singh et al., 2021) and UFold (Fu et al., 2022), which enhance accuracy by using neural networks and evolutionary data. Models such as E2Efold (Chen et al., 2020) and RNA-FM (Chen et al., 2022) employ transformer architectures to achieve state-of-the-art results in secondary structure prediction.
RNA 3D structure prediction has progressed through ab initio, template-based, and deep learning approaches. Ab initio methods (e.g., iFoldRNA (Sharma et al., 2008), SimRNA (Boniecki et al., 2016)) balance detail and efficiency but struggle with non-canonical interactions. Template-based models (e.g., FARNA/FARFAR (Das & Baker, 2007), 3dRNA (Zhang et al., 2020)) depend on existing structures but are limited by available data. Deep learning models like DeepFoldRNA (Pearce et al., 2022), RhoFold (Shen et al., 2022), RoseTTAFoldNA (Baek et al., 2024), and trRosettaRNA (Wang et al., 2023) show promise in predicting 3D structures from sequence data but face challenges with novel RNA families due to RNA’s conformational flexibility (Kulkarni et al., 2023).
Appendix B Models overview